The electric potential in a certain reasons of space along +x axis varies as Vx=3x2+5 where x is in meters and v is in volts. The electric field at point (4,0) is
 

Grade:

1 Answers

Arun
4 years ago
 
Electric potential V = 3x2+ 5 volts
The electric potential changes only along x-axis, We know that Ex = -dV/dx ? Ex = - d(6x)/ dx = - 8x
The electric field at point (4,0) will be (here x = 4) Ex – 24 volt